The price of admission for a ticket was $10 for adults and $3 for children. 520 tickets were sold, the resulting revenue was $3450. How many adults and how many children attended the concert?
Answer by checkley77(12569)

a+c=520 or a=520-c
10a+3c=3450
10(520-c)+3c=3450
5200-10c+3c=3450
-7c=3450-5200
-7c=-1750
c=1750/7
c=250 number of children tickets sold.
a+250=520
a=520-250
a=270 adult tickets sold.
Proof:
10*270+3*250=3450
2700+750=3450
3450=3450